# A Comprehensive Survey of Contamination Detection Methods in Large Language Models

### v4

@@ 4.3 Future Challenges
Real-time data contamination detection systems that can monitor data streams continuously and alert users to potential contamination events as they happen would be of upmost importance in critical applications like finance (Frésard et al., 2011) and healthcare (Gao et al., 2022) where data integrity is paramount and model reliability critical. Given the sheer volume of new data uploaded daily on the Internet, solving this task requires major technological breakthroughs.
@@
Dekoninck et al. (2024) show a very effective way to bypass several existing contamination detection methods: Evasive Augmentation Learning (EAL) consists in paraphrasing benchmarks with GPT-4 and fine-tuning the LLM on the paraphrased data. This method can bypass even oracle-access contamination detection methods, calling for the development of more robust methods.
@@
Ethical and legal frameworks (Chang, 2021) are needed to govern the collection, usage, and management of LLM pre-training data. This includes protocols for data privacy, consent, and use that help prevent the incorporation of contaminated data from unethical sources and the contamination of widely used pre-training data sources (e.g., CommonCrawl). Developing contamination detection techniques that do not compromise individual privacy will be crucial (Lin et al., 2010; Hayes & Ohrimenko, 2018).
@@ 5 Conclusion
As LLMs rapidly evolve and the training data become more extensive, LLMs’ performances are inevitably biased by contamination. Eventually, most (if not all) target-answer-based public evaluation datasets end up in commonly used pre-training data dumps, which are updated every few months. Therefore, to help detect and quantify contamination, in this survey, we investigated, structured, and classified the current landscape of contamination detection in LLMs. Depending on a researcher’s level of access to the data and LLM parameters, our survey can efficiently guide toward a set of relevant tools to detect contamination.

### v5

@@ 5.4 Future Challenges
Real-time data contamination detection systems that can monitor data streams continuously and alert users to potential contamination events as they happen would be of upmost importance in critical applications like finance (Frésard et al., 2011) and healthcare (Gao et al., 2022) where data integrity is paramount and model reliability critical. Given the sheer volume of new data uploaded daily on the Internet, solving this task requires major technological breakthroughs. The key difficulty lies in achieving high recall, to ensure that no contaminated data sample slips through undetected. We hypothesize that compression (Han et al., 2015), distillation (Hinton et al., 2015; Sanh et al., 2019) and quantization (Banner et al., 2018; Sun et al., 2020) techniques could play a key role in solving this problem, by decreasing model size, accelerating inference and scaling edge-device deployment of anomaly detection models or closed-data contamination detection techniques. Besides, leveraging federated learning and edge computing could enable decentralized, privacy-preserving detection at the source of data generation.
@@
Dekoninck et al. (2024) show a very effective way to bypass several existing contamination detection methods: Evasive Augmentation Learning (EAL) consists in paraphrasing benchmarks with GPT-4 and fine-tuning the LLM on the paraphrased data. This method can bypass even oracle-access contamination detection methods, calling for the development of more robust methods. An intuitive future line of defense to such techniques is to incorporate large-scale adversarial training (Jia & Liang, 2017; Liu et al., 2020) into contamination detection models, and train these models on benchmarks which have been paraphrased or noised. We expect to witness a cat-and-mouse game between techniques such as EAL, which contaminate under the radar, and contamination detection techniques catching up to the finest transformations of benchmarks, such as paraphrases.
@@
Ethical and legal frameworks (Chang, 2021) are needed to govern the collection, usage, and management of LLM pre-training data. A strong legal framework will help prevent the inclusion of data from unethical sources or from copyrighted sources, as well as help prevent the contamination of evaluation benchmarks into widely used pre-training data sources (e.g., CommonCrawl). However, adopting such a worldwide pre-training data legal framework would be challenging due to several reasons. First, the lines are blurry between public and private data, and many web-crawled datasets include a mix of public, open-access and proprietary data. Jurisdiction surrounding data varies from a country to another, complicating the adoption of global standards. Besides, the authorization to train machine learning models on copyrighted data remains unclear in many jurisdictions to this day. Another major blocker is the fact that pre-training data is nowadays treated with secrecy by large technological companies releasing flagship LLMs, as publishing the exact pre-training data would enable their competitors to catch up with their own LLMs. Lastly, enforcing and auditing a legal framework surrounding pre-training pipelines in practice would be very complex and costly. Above all, legal frameworks on pre-training data are needed to protect personal data privacy. The General Data Protection Regulation (GDPR), acted in 2018 in the European Union (EU), imposes strict rules on how organizations can collect, store and use personal data from EU citizens. With the exponential growth of pre-training corpora, it becomes harder and harder yet critical to ensure that personal data does not get included into LLM training sets. Once an LLM has seen and memorized personal data, there is no clear mechanism for an individual to exercise their right to erase their personal data from the LLM’s knowledge. In the near future, developing contamination detection techniques that do not compromise individual privacy will be crucial (Lin et al., 2010; Hayes & Ohrimenko, 2018).
@@ 6 Conclusion
As LLMs rapidly evolve and the training data become more extensive, LLMs’ performances are inevitably biased by contamination. Eventually, most (if not all) target-answer-based public evaluation datasets end up in commonly used pre-training data dumps, which are updated every few months. Therefore, to help detect and quantify contamination, in this survey, we investigated, structured, and classified the current landscape of contamination detection in LLMs. Depending on a researcher’s level of access to the data and to the LLM parameters, our survey can quickly guide them toward a set of relevant tools to detect contamination.